# On computability of equilibrium states

### v2

@@ 5. Hyperbolic rational maps
Next, we verify that $B_{d_{\widehat{\mathbb{C}}}}(f(x),\xi)\subseteq f(B_{d_{\widehat{\mathbb{C}}}}(x,\eta))$. To see this, since $V_{4}\leqslant\frac{d_{\widehat{\mathbb{C}}}(x,y)}{\lvert x-y\rvert}\leqslant 2$, it suffices to check that $B(f(x),2\xi)\subseteq f(B(x,l))$. By statement (i), it follows from by Köbe’s distorsion theorem that
$B(f(x),2\xi)=B(f(x),V_{2}l/4)\subseteq B(f(x),\lvert f^{\prime}(x)\rvert l/4)\subseteq f(B(x,l)).$
This completes the proof of statement (ii). ∎
@@
Let $-^f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Let $x_{k}$, $1\leqslant k\leqslant m$, be an enumeration of $f^{-1}(\infty)\cup\{\infty\}$, and $r>0$ be a constant satisfying that $\mathcal{J}_{f}\subseteq K\coloneqq\bigcap_{k=1}^{m}B_{d_{\widehat{\mathbb{C}}}}(x_{k},r)^{c}$ and $B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)\cap B_{d_{\widehat{\mathbb{C}}}}(x_{j},r)=\emptyset$ for all $1\leqslant i<j\leqslant m$. Fix constants $D_{1},\,D_{2},\,C_{1},\,C_{2},\,C>0$ satisfy the following conditions:
@@
    $\lvert f^{\#}(z)\rvert\geqslant C$ for each $z\in\mathcal{J}_{f}$.
Then the following statements hold:
@@
    $d_{K}(x,y)\coloneqq\inf\biggl{\{}\int_{\gamma}\!\mathrm{d}s:\gamma(0)=x,\,\gamma(1)=y,\,\text{and }\gamma\colon[0,1]\rightarrow K\text{ is continuous}\biggr{\}};$
@@
    If $g\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a rational map, then $\lvert g\rvert$ is $\bigl{(}A(1+D_{1}^{2})\pi/2\bigr{)}$-Lipschitz continuous with respect to $d_{\widehat{\mathbb{C}}}$ on $K$, where $A\coloneqq\sup_{z\in K}\lvert g^{\prime}(z)\rvert$;
@@
    $\varphi_{f}(z)$ is $(BC^{-1})$-Lipschitz continuous on $\mathcal{J}_{f}$ with respect to $d_{\widehat{\mathbb{C}}}$, where $\varphi_{f}(z)\coloneqq\log\bigl{(}f^{\#}(z)\bigr{)}$ for each $z\in\mathcal{J}_{f}$, where $B\coloneqq(1+D_{1}^{2})\bigl{(}C_{2}(1+D_{1}^{2})+D_{2}C_{1}^{2}(1+D_{1}^{2})+C_{1}D_{1}\bigr{)}\pi$.
@@
(i) The path-connectedness of $K$ is immediate from its definition. Hence, $d_{K}$ is well-defined. By subdividing the spherical geodesic from $x$ to $y$, we only need to consider the case where $x,\,y\in\partial B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)$ for some $1\leqslant i\leqslant m$. Then we have
$d_{K}(x,y)\leqslant\pi\sigma(x,y)\leqslant\pi d_{\widehat{\mathbb{C}}}(x,y).$
(ii) Fix an arbitrary pair of $x,\,y\in K$. Since $K$ is closed, there exists a continuous function $\gamma\colon[0,1]\mapsto K$ with $\gamma(0)=x,\,\gamma(1)=y,$ and $d_{K}(x,y)=\int_{\gamma}\!\mathrm{d}s$. Hence, for each pair of $x,\,y\in K$, we have
$\lvert g(x)-g(y)\rvert=\biggl{\lvert}\int_{\gamma}\!g^{\prime}(z)\,\mathrm{d}z\biggr{\rvert}\leqslant\int_{\gamma}\!\frac{A(1+\lvert z\rvert^{2})}{2}\,\mathrm{d}s(z)\leqslant\frac{A(1+D_{1}^{2})d_{K}(x,y)}{2}\leqslant\frac{A(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y)}{2}.$
(iii) Define $u(x)=1+\lvert x\rvert^{2}$, $v(x)\coloneqq\frac{1}{1+\lvert f(x)\rvert^{2}}$, and $w(x)\coloneqq\lvert f^{\prime}(x)\rvert$ for each $x\in K$. Then we establish that $u,v,$ and $w$ are all Lipschitz continuous functions, respectively. By statement (i), conditions (1) and (2), we have
(5.3) $\lvert u(x)-u(y)\rvert=(\lvert x\rvert+\lvert y\rvert)\left\lvert\lvert x\rvert-\lvert y\rvert\right\rvert\leqslant 2D_{1}\lvert x-y\rvert\leqslant D_{1}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y),$
(5.4) $\lvert v(x)-v(y)\rvert=\frac{\lvert\lvert f(x)\rvert^{2}-\lvert f(y)\rvert^{2}\rvert}{(1+\lvert f(x)\rvert^{2})(1+\lvert f(y)\rvert^{2})}\leqslant 2D_{2}\lvert\lvert f(x)\rvert-\lvert f(y)\rvert\rvert\leqslant D_{2}C_{1}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y),$
@@
(5.5) $\lvert w(x)-w(y)\rvert\leqslant C_{2}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in K$. By the definition of $K$, (5.1) is true for each $z\in K$. Moreover, by (5.3), (5.4), and (5.5), we have
(5.6) $\bigl{\lvert}f^{\#}(x)-f^{\#}(y)\bigr{\rvert}$
$\qquad\leqslant\lvert u(x)v(x)\rvert\lvert w(x)-w(y)\rvert+\lvert u(x)w(y)\rvert\lvert v(x)-v(y)\rvert+\lvert v(y)w(y)\rvert\lvert u(x)-u(y)\rvert$
$\qquad\leqslant\bigl{(}C_{2}(1+D_{1}^{2})^{2}+D_{2}C_{1}^{2}(1+D_{1}^{2})^{2}+C_{1}D_{1}(1+D_{1}^{2})\bigr{)}\pi d_{\widehat{\mathbb{C}}}(x,y)$
$\qquad=Bd_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in K$. Note that $\log x<x-1$ for each $x>1$. Therefore, it follow from condition (3) that
@@
Now we are ready to establish Theorem C.
@@
Since $f$ is a hyperbolic rational map, it follows from Lemma 5.4 that there exists a computable point $u$ in $\mathcal{J}_{f}^{c}$. By Lemma 5.5, we can compute a point $u\in\mathcal{J}_{f}^{c}$. Define a new function $g$ by
@@
$g(z)\coloneqq\bigl{(}U\circ f\circ U^{-1}\bigr{)}(z)=1\big{/}\bigl{(}f\bigl{(}u+z^{-1}\bigr{)}-u\bigr{)}$
for each $z\in\widehat{\mathbb{C}}$. It follows from the computability of $f$ and $u$ that $g$ is a computable hyperbolic rational map with $\infty\notin\mathcal{J}_{g}$. First, we compute constants $n,\,\eta,\,\lambda$, and $\xi$ which satisfy that $g^{n}$ is distance-expanding on $\mathcal{J}_{g}$ with respect to the spherical metric with $\eta,\,\lambda$, and $\xi$. We apply a standard root-finding algorithm to compute the periodic points of $g$ as well as their multipliers. We can semi-decide whether a periodic point is repelling (if the multiplier is greater than $1$). Hence, we can compute the repelling periodic point $w$ of $g$. By Proposition 5.6, one can see that $\mathcal{J}_{g}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{g},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}g^{-n}(w)\bigr{)}$. By [BRY12, Proposition 2.13], we can compute a constant $r_{0}>0$ with $\lvert z\rvert\leqslant r_{0}$ for each $z\in\mathcal{J}_{g}$. Then by (5.2), one can conclude that
@@
$d_{\widehat{\mathbb{C}}}(x,y)\geqslant\sigma(x,y)\geqslant 2\lvert x-y\rvert\big{/}\bigl{(}1+r_{0}^{2}\bigr{)}$
[…]

### v3

@@ 5. Hyperbolic rational maps
Next, we verify that $B_{d_{\widehat{\mathbb{C}}}}(f(x),\xi)\cap\mathcal{J}_{f}\subseteq f\bigl{(}B_{d_{\widehat{\mathbb{C}}}}(x,\eta)\cap\mathcal{J}_{f}\bigr{)}$ for each $x\in\mathcal{J}_{f}$. Indeed, by $d_{\widehat{\mathbb{C}}}(x,y)\leqslant 2\lvert x-y\rvert$ for each pair of $x,\,y\in\widehat{\mathbb{C}}$, we have
(5.3) $B_{d_{\widehat{\mathbb{C}}}}(f(x),\xi)\subseteq B(f(x),2\xi)\quad\text{ for each }x\in\widehat{\mathbb{C}}$.
By the definitions of $l,\,\xi$, property (2), statement (i), and Koebe’s one-quarter theorem, we have
(5.4) $B(f(x),2\xi)=B(f(x),V_{2}l/8)\subseteq B(f(x),\lvert f^{\prime}(x)\rvert l/8)\subseteq f(B(x,l/2))\quad\text{ for each }x\in\mathcal{J}_{f}.$
By property (3), we have $B(x,l/2)\cap\mathcal{J}_{f}\subseteq B_{d_{\widehat{\mathbb{C}}}}(x,\eta)\cap\mathcal{J}_{f}$. Note that by [CG93, Theorem III.1.3], $f(\mathcal{J}_{f})=\mathcal{J}_{f}$. Then $f(A\cap\mathcal{J}_{f})=f(A)\cap\mathcal{J}_{f}$ for each subset $A\subseteq\widehat{\mathbb{C}}$. Thus, we obtain that
(5.5) $f(B(x,l/2))\cap\mathcal{J}_{f}=f(B(x,l/2)\cap\mathcal{J}_{f})\subseteq f\bigl{(}B_{d_{\widehat{\mathbb{C}}}}(x,\eta)\cap\mathcal{J}_{f}\bigr{)}\quad\text{ for each }x\in\mathcal{J}_{f}.$
Therefore, by (5.3), (5.4), and (5.5), we have $B_{d_{\widehat{\mathbb{C}}}}(f(x),\xi)\cap\mathcal{J}_{f}\subseteq f\bigl{(}B_{d_{\widehat{\mathbb{C}}}}(x,\eta)\cap\mathcal{J}_{f}\bigr{)}$ for each $x\in\mathcal{J}_{f}$. This completes the proof of statement (ii). ∎
@@
Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be a hyperbolic rational map of degree $d\geqslant 2$ with $\infty\notin\mathcal{J}_{f}$. Let $x_{k}$, $1\leqslant k\leqslant m$, be an enumeration of $f^{-1}(\infty)\cup\{\infty\}$, and $r>0$ be a constant satisfying $\mathcal{J}_{f}\subseteq K\coloneqq\bigcap_{k=1}^{m}B_{d_{\widehat{\mathbb{C}}}}(x_{k},r)^{c}$ and $B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)\cap B_{d_{\widehat{\mathbb{C}}}}(x_{j},r)=\emptyset$ for all $1\leqslant i<j\leqslant m$. Assume that there exist positive constants $D_{1},\,D_{2},\,C_{1},\,C_{2}$, and $C$ satisfying the following properties:
@@
    $\left\lvert f^{\#}(z)\right\rvert\geqslant C$ for each $z\in\mathcal{J}_{f}$.
Then the following statements are true:
@@
    $d_{K}(x,y)\coloneqq\inf\biggl{\{}\int_{\gamma}\mathrm{d}s:\gamma(0)=x,\,\gamma(1)=y,\text{ and }\gamma\colon[0,1]\rightarrow K\text{ is continuous}\biggr{\}}.$
@@
    If $g\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ is a rational map, then the restriction $\lvert g\rvert|_{K}\colon K\rightarrow\mathbb{R}$ of the norm of $g$ is $\bigl{(}A\bigl{(}1+D_{1}^{2}\bigr{)}\pi/2\bigr{)}$-Lipschitz continuous with respect to $d_{\widehat{\mathbb{C}}}$, where $A\coloneqq\sup_{z\in K}\lvert g^{\prime}(z)\rvert$.
@@
    The restriction $\varphi_{f}|_{\mathcal{J}_{f}}\colon\mathcal{J}_{f}\rightarrow\mathbb{R}$ is $\bigl{(}BC^{-1}\bigr{)}$-Lipschitz continuous with respect to $d_{\widehat{\mathbb{C}}}$, where $\varphi_{f}(z)\coloneqq\log\bigl{(}f^{\#}(z)\bigr{)}$ for each $z\in\mathcal{J}_{f}$ and $B\coloneqq\bigl{(}1+D_{1}^{2}\bigr{)}\bigl{(}C_{2}\bigl{(}1+D_{1}^{2}\bigr{)}+D_{2}C_{1}^{2}\bigl{(}1+D_{1}^{2}\bigr{)}+C_{1}D_{1}\bigr{)}\pi$.
@@
(i) The path-connectedness of $K$ is immediate from its definition. Hence, $d_{K}$ is well-defined. Consider an arbitrary pair of $x,\,y\in K$. By subdividing the spherical geodesic from $x$ to $y$, we only need to consider the case where $x,\,y\in\partial B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)$ for some integer $1\leqslant i\leqslant m$. Then we have $d_{K}(x,y)\leqslant\pi\sigma(x,y)\leqslant\pi d_{\widehat{\mathbb{C}}}(x,y).$
(ii) Fix an arbitrary pair of $x,\,y\in K$. By the definition of $K$, $K$ is closed. Then there exists a continuous function $\gamma\colon[0,1]\rightarrow K$ with $\gamma(0)=x,\,\gamma(1)=y,$ and $d_{K}(x,y)=\int_{\gamma}\mathrm{d}s$. Hence, we have
$\lvert\lvert g(x)\rvert-\lvert g(y)\rvert\rvert\leqslant\biggl{\lvert}\int_{\gamma}\!g^{\prime}(z)\,\mathrm{d}z\biggr{\rvert}\leqslant\int_{\gamma}\!\frac{A(1+\lvert z\rvert^{2})}{2}\,\mathrm{d}s(z)\leqslant\frac{A\bigl{(}1+D_{1}^{2}\bigr{)}d_{K}(x,y)}{2}\leqslant\frac{A\bigl{(}1+D_{1}^{2}\bigr{)}\pi d_{\widehat{\mathbb{C}}}(x,y)}{2}.$
This completes the proof of statement (ii).
(iii) Define $u(x)\coloneqq 1+\lvert x\rvert^{2}$, $v(x)\coloneqq\frac{1}{1+\lvert f(x)\rvert^{2}}$, and $w(x)\coloneqq\lvert f^{\prime}(x)\rvert$ for each $x\in K$.
By statement (ii), conditions (1) and (2), we have
(5.6) $\lvert u(x)-u(y)\rvert=(\lvert x\rvert+\lvert y\rvert)\left\lvert\lvert x\rvert-\lvert y\rvert\right\rvert\leqslant 2D_{1}\left\lvert\lvert x\rvert-\lvert y\rvert\right\rvert\leqslant D_{1}(1+D_{1}^{2})\pi d_{\widehat{\mathbb{C}}}(x,y),$
(5.7) $\lvert v(x)-v(y)\rvert=\frac{\lvert\lvert f(x)\rvert^{2}-\lvert f(y)\rvert^{2}\rvert}{(1+\lvert f(x)\rvert^{2})(1+\lvert f(y)\rvert^{2})}\leqslant 2D_{2}\lvert\lvert f(x)\rvert-\lvert f(y)\rvert\rvert\leqslant D_{2}C_{1}\bigl{(}1+D_{1}^{2}\bigr{)}\pi d_{\widehat{\mathbb{C}}}(x,y),$
@@
(5.8) $\lvert w(x)-w(y)\rvert\leqslant C_{2}\bigl{(}1+D_{1}^{2}\bigr{)}\pi d_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in K$. By the definition of $K$, (5.1) holds for each $z\in K$. Moreover, by (5.6), (5.7), and (5.8), we have
[…]
$\leqslant\bigl{(}C_{2}\bigl{(}1+D_{1}^{2}\bigr{)}^{2}+D_{2}C_{1}^{2}\bigl{(}1+D_{1}^{2}\bigr{)}^{2}+C_{1}D_{1}\bigl{(}1+D_{1}^{2}\bigr{)}\bigr{)}\pi d_{\widehat{\mathbb{C}}}(x,y)=Bd_{\widehat{\mathbb{C}}}(x,y)$
for each pair of $x,\,y\in K$. Note that $\log x\leqslant x-1$ for each $x\geqslant 1$. Then it follows from condition (3) and (5.9) that
@@
Now, we are ready to establish Theorem C.
@@
Since $f$ is a hyperbolic rational map, it follows from Lemma 5.4 and the compactness of $\mathcal{J}_{f}$ that there exists a computable point $u$ in $\mathcal{J}_{f}^{c}\cap\mathbb{Q}^{2}$. By Lemma 5.5, there exists an algorithm which on input $u\in\mathbb{Q}^{2}$, halts if and only if $d_{\widehat{\mathbb{C}}}(u,\mathcal{J}_{f})>0$. With this algorithm, we can compute a point $u\in\mathcal{J}_{f}^{c}$. Define, for each $z\in\widehat{\mathbb{C}}$,
@@
$U(z)\coloneqq\frac{\overline{u}z+1}{u-z}\quad\text{ and }\quad g(z)\coloneqq\bigl{(}U\circ f\circ U^{-1}\bigr{)}(z).$
Then $U$ is a Möbius transformation. Since the function $f$ and the point $u$ are both computable, by the definition of $g$, $g$ is also a computable hyperbolic rational map with $\mathcal{J}_{g}=U(\mathcal{J}_{f})$. Since $u\in\mathcal{J}_{f}^{c}$, we have $\infty\in\mathcal{J}_{g}^{c}$.
By [CG93, Theorem III.1.2], $\mathcal{J}_{g}\neq\emptyset$. Then by [CG93, Theorem III.3.1], there exists a repelling periodic point of $g$, say $w$. By Proposition 5.6, $\bigl{(}\mathcal{J}_{g},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}g^{-n}(w)\bigr{)}$ is a computable metric space and $\mathcal{J}_{g}$ is recursively compact in the computable metric space $\bigl{(}\mathcal{J}_{g},\,d_{\widehat{\mathbb{C}}},\,\bigcup_{n\in\mathbb{N}}g^{-n}(w)\bigr{)}$.
Now we apply Proposition 5.7 to compute constants $n\in\mathbb{N},\,\eta>0,\,\lambda>1$, and $\xi>0$ satisfying that the restriction $g^{n}|_{\mathcal{J}_{g}}$ is distance-expanding with respect to the spherical metric with constants $\eta,\,\lambda$, and $\xi$ above.
Put $V_{4}\coloneqq\frac{2}{1+r_{0}^{2}}$. By [BRY12, Proposition 2.13], we can compute a constant $r_{0}>0$ with $\lvert z\rvert\leqslant r_{0}$ for each $z\in\mathcal{J}_{g}$. Then by (5.2), one can conclude that
@@
$d_{\widehat{\mathbb{C}}}(x,y)\geqslant\sigma(x,y)\geqslant 2\lvert x-y\rvert\big{/}\bigl{(}1+r_{0}^{2}\bigr{)}=V_{4}\lvert x-y\rvert\quad\text{ for each pair of }x,\,y\in\mathcal{J}_{g}.$
By Lemma 5.3, there exists sufficiently large $n\in\mathbb{N}$ with $\lvert(g^{n})^{\prime}(z)\rvert>4(1+r_{0})^{2}$ for each $z\in\mathcal{J}_{g}$. Note that by [BRY12, Proposition 2.13], there exists an algorithm which on input $n\in\mathbb{N}$, computes $\inf_{z\in\mathcal{J}_{g}}\lvert(g^{n})^{\prime}(z)\rvert$. Hence, we can apply this algorithm to compute three constants $n\in\mathbb{N},\,V_{2}>0$, and $V_{3}>0$ satisfying
$V_{3}\geqslant\lvert(g^{n})^{\prime}(z)\rvert\geqslant V_{2}>4(1+r_{0})^{2}\quad\text{ for each }z\in\mathcal{J}_{g}.$
Since $g$ is computable, we can apply a standard root-finding algorithm (see e.g. [BBY12, Appendix A]) to compute all the points in $\bigl{\{}z\in\widehat{\mathbb{C}}:(g^{n})^{\prime}(z)=0\bigr{\}}$. Hence, we can compute all the points in the set $C_{g^{n}}$ of the critical values of $g^{n}$. By Lemma 5.5, we can compute a constant $V_{1}>0$ satisfying that $d(z,\mathcal{J}_{g})\geqslant V_{1}$ for each $z\in C_{g^{n}}$. Note that, by [CG93, Theorem III.1.4], $\mathcal{J}_{g^{n}}=\mathcal{J}_{g}$. Then it is not hard to see that the above constants $V_{1},\,V_{2},\,V_{3},\,V_{4}>0$ satisfy Properties (1), (2), and (3) in Proposition 5.7 for the hyperbolic rational map $g^{n}$. Hence, by Proposition 5.7, we can compute the constants $\eta$, $\lambda,$ and $\xi$ satisfying that the restriction $g^{n}|_{\mathcal{J}_{g}}$ is distance-expanding with respect to the spherical metric with constants $\eta,\,\lambda$, and $\xi$.
Next, we establish that $\varphi_{g^{n}}$ is $L$-Lipschitz continuous for some $L>0$ and compute one of such constant $L$. Since $g^{n}$ is a computable rational map, we can compute $(g^{n})^{-1}(\infty)\cup\{\infty\}$, and we enumerate the points as $x_{k},\,1\leqslant k\leqslant m$. Then by Lemma 5.5, we can compute $r>0$ satisfying $\mathcal{J}_{g}\subseteq K\coloneqq\bigcap_{k=1}^{m}B_{d_{\widehat{\mathbb{C}}}}(x_{k},r)^{c}$ and $B_{d_{\widehat{\mathbb{C}}}}(x_{i},r)\cap B_{d_{\widehat{\mathbb{C}}}}(x_{j},r)=\emptyset$ for all $1\leqslant i<j\leqslant m$. By Proposition 3.15 (iii), $K$ is recursively compact. Hence, by Proposition 3.15 (iv), we can compute $D_{1}$, $D_{2}$, $C_{1}$, $C_{2},\,C>0$ satisfying conditions (1), (2), and (3) of Proposition 5.8 for the map $g^{n}$. Hence, by Proposition 5.8, we can compute $L>0$ satisfying that $\varphi_{g^{n}}$ is $L$-Lipschitz continuous with respect to the spherical metric.
We are now ready to return to the proof of the main statement. By [CG93, Theorem III.3.2], $g^{n}|_{\mathcal{J}_{g}}$ is topologically exact. Moreover, by Propositions 5.65.7, and 5.8, for each repelling periodic point $w$ of $g$, the septet $\bigl{(}\widehat{\mathbb{C}},\,\mathcal{J}_{g},\,d_{\widehat{\mathbb{C}}},\,\mathbb{Q}^{2},\,\bigcup_{n\in\mathbb{N}}g^{-n}(w),\,t\varphi_{g^{n}},\,g^{n}\bigr{)}$ satisfies the Assumption B in Section 1. Moreover, by the above analysis, we can compute the five constants in Theorem 1.2 (iii) and (iv) for the map $g^{n}$ and the function $t\varphi_{g^{n}}$. Proposition 5.6 (ii) gives the algorithm in Theorem 1.2 (i). Hence, by Theorem 1.2, the unique equilibrium state for $g^{n}$ and $t\varphi_{g^{n}}$ is computable.
By [PU10, Theorem 2.4.6 (a) and Lemma 3.2.7], the unique equilibrium state for $g$ and $t\varphi_{g}$ is exactly the unique equilibrium state for $g^{n}$ and $tS_{n}\varphi_{g}$. Moreover, by (5.1), $S_{n}\varphi_{g}=\varphi_{g^{n}}$, because $\infty\notin\mathcal{J}_{g}$. Hence, the unique equilibrium state for $g$ and $t\varphi_{g}$ is computable. By simple computation, for each $z\in\widehat{\mathbb{C}}$,
$\mathrm{d}s(U(z))=\frac{2\lvert\mathrm{d}(U(z))\rvert}{1+\lvert U(z)\rvert^{2}}=\frac{2\lvert\mathrm{d}z\rvert\lvert U^{\prime}(z)\rvert}{1+\lvert U(z)\rvert^{2}}=\frac{2\lvert\mathrm{d}z\rvert(1+\lvert u\rvert^{2})}{\lvert u-z\rvert^{2}+\lvert\overline{u}z+1\rvert^{2}}=\frac{2\lvert\mathrm{d}z\rvert}{1+\lvert z\rvert^{2}}=\mathrm{d}s(z).$
Hence, the map $U$ preserves the spherical metric. Therefore, the unique equilibrium state for $f$ and $t\varphi_{f}$ is also computable. ∎